OpenStudy (anonymous):

PLEASE Show all work SO I COULD UNDERSTAND. Find the following quotient: 9X^2-25Y^2/3X+5Y

OpenStudy (mertsj):

\[\frac{9x^2-25y^2}{3x+5y}=\frac{(3x-5y)(3x+5y)}{3x+5y}=3x-5y\]

OpenStudy (mertsj):

Notice that you cancel out the common factor 3x+5y

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Do you understand?

OpenStudy (anonymous):

not really am trying to see how you did your problem

OpenStudy (mertsj):

The numerator is the difference of two squares so I used the property that: \[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

and that how u got (3x-5y)(3x+5y)(3x+5y)=3x-5y

OpenStudy (anonymous):

help i need to see the work or i cant understand

OpenStudy (mertsj):